1. A weighted coin has a 40% probability of coming up heads.

g. What is the probability there will be more than 35 heads in 100 flips?
Answer #1

X ~ Binomial (n,p)

Where b = 100, p = 0.40

Mean = np = 100 * 0.40 = 40

Standard deviation = Sqrt (np(1-p))

= sqrt(100 * 0.40 * 0.60)

= 4.8990

Using normal approximation,

P( X > 35) = p( Z > x+0.5 - mean / SD)

= P( Z > 35.5 - 40 / 4.8990)

= P( Z > -0.9186)

= P( Z < 0.9186)

= 0.8208